# Improved QLDPC Surgery: Logical Measurements and Bridging Codes

### v3

@@ 3.8 Bridging Codes
This limitation is fully resolved in [SJOY24], where it was shown that when $F^{(1)}$, $F^{(2)}$ are simple expander graphs (such as those constructed in [WY24, IGND24]), one can always efficiently solve Problem 26 to guarantee a LDPC bridge construction. This means the adapters constructed in [SJOY24] are truly universal, in the sense that they can connect arbitrary QLDPC code families while maintaining fault-tolerance and LDPC. We refer readers to [SJOY24] for further studies and applications of the idea of code adapting.
@@ 4 Case Study: Gross Code
Since the polynomials are over $\mathbb{F}_{2}$, they are also naturally interpreted as subsets of $\mathcal{M}$. Then the Tanner graph is given as follows. For some $m\in\mathcal{M}$, the vertex $X(m)$ is connected to the vertices $L(Am):=\{L(m^{\prime})$ for $m^{\prime}$ in $Am\}$, and to $R(Bm):=\{R(m^{\prime\prime})$ for $m^{\prime\prime}$ in $Bm\}$. Similarly, if $A^{\top}$ is the polynomial consisting of the inverses of the monomials of $A$, that is, $A^{T}=x^{-3}+y^{-1}+y^{-2}$, then the vertex $Z(m)$ is connected to $L(B^{\top}m):=\{L(m^{\prime})$ for $m^{\prime}$ in $B^{\top}m\}$, and to $R(A^{\top}m):=\{R(m^{\prime\prime})$ for $m^{\prime\prime}$ in $A^{\top}m\}$.
@@ 4.2 Numerical simulations
We perform three experiments on the syndrome circuits. First, we perform a memory experiment to emperically determine the circuit-level distance and to compare to the code’s idle cycle. Second, we determine how many syndrome cycles are suitable for a logical measurement by comparing the logical error rate to the measurement error rate. Both of these experiments rely on the modular decoding approach described in Section 3.5. In the third experiment we assess the benefits of the modular decoding approach over a simpler implementation relying on BPOSD alone. Our software is implemented in python and leverages the stim [Gid21], pymatching [HG23], and bp_osd [RWBC20, Rof22] libraries.
@@ 4.3 Implementing logical Clifford gates
There are some variations of (70). The |$\ket{0}$ initialization can be performed through a $Z$ measurement, resulting in a $Z\to X\to Y$ sequence of basis changes. By inserting single-qubit Clifford gates and their inverses, this sequence can be changed to any ordering of the $X,Y,Z$ bases.
@@ 5 Future Work
Ancilla systems are particularly affordable when only one layer of additional qubits is necessary. We were able to find logical operators on the gross code where this is the case. Are there any other families of codes where this is possible? Our analytical techniques did not suffice to prove that a monolayer system suffices for the gross code. Is it ever possible to prove that a single layer is enough? Quantum expander codes seem promising candidates.
Measurements using ancilla systems allow us to implement Clifford gates through techniques similar to lattice surgery [Lit19]. Can these methods be extended to inject physical magic states into a logical qubit? Such an operation could form the foundation for a magic state factory on an LDPC code.
We derived explicit syndrome circuits for the $\bar{X}$ and $\bar{Z}$ measurement schemes for the gross code. Are these syndrome circuits optimal in terms of the number of fault locations? Can one design optimal or otherwise high-performance syndrome circuits for other measurement schemes, perhaps assuming that the original code $\mathcal{G}$ also has an efficient syndrome circuit?
From Section 4.3 it is clear that implementing logical computations will require very many measurements, which in turn required several rounds of error correction. For certain error correction codes, such as perhaps [SHR24], is there a single-shot error correction scheme for merged codes?
@@ A.1 Proofs for the $X$ Ancilla System
Finally, we show that if $\mathcal{G}$ has $k$ logical qubits, then $\mathcal{G}_{X}$ has $k-1$ logical qubits. We proceed by induction on the number of layers, and we use $\mathcal{G}^{l}_{X}$ to indicate the merged code for measuring $\bar{X}_{M}$ with $l$ layers, and $H^{X}_{l}$, $H^{Z}_{l}$ its parity check matrices.
@@
The size of $U_{1}$ is the rank of $G$. Observe that the rank of $F$ is $|V|-1$, since we assumed that the only logical operator supported on $V$ is $\bar{X}_{M}$. Therefore $|U_{1}|=\operatorname{rank}(G)=|C_{0}|-\operatorname{rank}(F)=|C_{0}|-|V_{0}|+1$, and the number logical qubits in $\mathcal{G}_{X}^{1}$ is
@@
We proceed with the inductive step: assuming that $\mathcal{G}^{L-2}_{X}$ has $k-1$ logical qubits, we show the same for $\mathcal{G}^{L}_{X}$. The newly added set of $X$ checks $V_{L}$ is linearly independent. A product of any subset of $V_{L}$ will have non-trivial support on $C_{L}$, and while $H_{X}(V_{L})=X(V_{L-1})$ has no support on $C_{L}$, by assumption it cannot be a stabilizer in $H^{X}_{L-2}$. Similarly, the newly added $Z$ checks $C_{L-1}$ is linearly independent, and their products always have support on $C_{L}$. We conclude that $\mathcal{G}_{X}^{L}$ has $k-1$ logical qubits.
In fact, these $k-1$ qubits are the remaining logical qubits from $\mathcal{G}$. For any $\bar{Z}$ that commutes with $\bar{X}_{M}$, its support on $V_{0}$ is even in size, and as a consequence of Lemma 8, we can clean its support from $V_{0}$ by multiplying by $Z$ checks of $\mathcal{G}$. This implies some $\bar{Z}^{\prime}\equiv_{\mathcal{G}}\bar{Z}$ supported only on $R$, which is then also a logical operator of $\mathcal{G}_{X}$. For a logical operator $\bar{X}\not\equiv_{\mathcal{G}}\bar{X}_{M}$ of $\mathcal{G}$ supported only on $V_{0}\cup R$, notice that $\bar{X}$ is also a logical operator of $\mathcal{G}_{X}$. It is nontrivial in $\mathcal{G}_{X}$ because there was some $\bar{Z}$ that anticommuted with $\bar{X}$ in $\mathcal{G}$ and commuted with $\bar{X}_{M}$, and therefore the construction of $\bar{Z}^{\prime}$ still anticommutes with $\bar{X}$ in $\mathcal{G}_{X}$. ∎
@@ A.2 Proofs for the $Y$ Ancilla System
which implies the number of logical qubits is $k-1$.
Inductively, we make the same observation as in the proof of Theorem 1: the newly added set of $X$ checks $V^{X}_{L}$ is linearly independent. Any $H_{X}(v\in V^{X}_{L})$ will have non-trivial support on $C_{L}$, and while $H_{X}(V^{X}_{L})=X(V_{L-1})$ has no support on $C_{L}$, by assumption it cannot be a stabilizer in $\mathcal{G}^{L-2}_{Y}$. Similarly, the newly added $Z$ checks in $C^{X}_{L-1}$ are linearly independent, and their products always have support on $C_{L}$. We conclude that $\mathcal{G}^{L}_{Y}$ has $k-1$ logical qubits, which are the original logical qubits from $\mathcal{G}$. ∎
@@ A.3 Proofs for the Multi-block $XX$ Ancilla System
Let $\mathcal{S}$ be the stabilizer group of the original codes and each ancilla qubit, including bridge qubits, in |$\ket{0}$, and $\mathcal{S}_{XX}$ be the stabilizer group of the joint merged code, Figure 6. Also let $\mathcal{L}$ be the group of logical operators of the unmeasured $12k-$ 1 qubit logical space and =$\mathcal{L}^{*}=\mathcal{L}\setminus\{I\}$. The versions of Lemmas 9 and 10 that apply to the joint code imply that the measurement fault distance is $\min(R,d_{Z}(\bar{Z}_{1}\mathcal{L}\mathcal{S}))$ and the logical fault distance is $\min(d_{Z}(\mathcal{L}^{*}\mathcal{S}),d_{X}(\mathcal{L}^{*}\mathcal{S}_{XX}))$. The original code $\mathcal{S}$ has code distance at least $d$ – if it did not, then the individual systems would not have fault distances at least $d$ – and so $d_{Z}(\bar{Z}_{1}\mathcal{L}\mathcal{S})\geq d$ and $d_{Z}(\mathcal{L}^{*}\mathcal{S})\geq d$. To complete the proof, it remains only to show that the $X$-distance of the joint merged code is good, i.e. $d_{X}(\mathcal{L}^{*}\mathcal{S}_{XX})\geq d$.
@@
×$\times\left(\prod_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}X(v_{j}F^{(1)\top}\in C_{j}^{(1)})\right).$ (102)
Since $|X(B)S_{1}|$ is at least the weight of factors in line 101, we proceed by ignoring the factors in line 102.

### v4

@@ 3.8 Bridging Codes
This limitation is fully resolved in [SJOY24], where it was shown that when $F^{(1)}$, $F^{(2)}$ are simple expander graphs (such as those constructed in [WY24, IGND25]), one can always efficiently solve Problem 26 to guarantee a LDPC bridge construction. This means the adapters constructed in [SJOY24] are truly universal, in the sense that they can connect arbitrary QLDPC code families while maintaining fault-tolerance and LDPC. We refer readers to [SJOY24] for further studies and applications of the idea of code adapting.
@@ 4 Case Study: Gross Code
Since the polynomials are over $\mathbb{F}_{2}$, they are also naturally interpreted as subsets of $\mathcal{M}$. Then the Tanner graph is given as follows. For some $m\in\mathcal{M}$, the vertex $X(m)$ is connected to the vertices $L(Am):=\{L(m^{\prime})$ for $m^{\prime}$ in Am}Am\}, and to $R(Bm):=\{R(m^{\prime\prime})$ for $m^{\prime\prime}$ in Bm}Bm\}. Similarly, if $A^{\top}$ is the polynomial consisting of the inverses of the monomials of $A$, that is, $A^{T}=x^{-3}+y^{-1}+y^{-2}$, then the vertex $Z(m)$ is connected to $L(B^{\top}m):=\{L(m^{\prime})$ for $m^{\prime}$ in $B^{\top}m\}$, and to $R(A^{\top}m):=\{R(m^{\prime\prime})$ for $m^{\prime\prime}$ in $A^{\top}m\}$.
@@ 4.2 Numerical simulations
We perform three experiments on the syndrome circuits. First, we perform a memory experiment to emperically determine the circuit-level distance and to compare to the code’s idle cycle. Second, we determine how many syndrome cycles are suitable for a logical measurement by comparing the logical error rate to the measurement error rate. Both of these experiments rely on the modular decoding approach described in Section 3.5. In the third experiment we assess the benefits of the modular decoding approach over a simpler implementation relying on BPOSD alone. Our software is implemented in python and leverages the stim [Gid21], pymatching [HG25], and bp_osd [RWBC20, Rof22] libraries.
@@ 4.3 Implementing logical Clifford gates
[…]
@@ 5 Future Work
In this work, we integrated three new ideas into the study of QLDPC surgery: using graph expansion to lower space overhead, gauge-fixing extraneous logical qubits to enhance fault-tolerance, and building bridge systems to connect multiple ancilla systems and codes. As demonstrated on the gross code, the gauge-fixed surgery scheme is a promising method of performing low-overhead logical computation on QLDPC codes. We sample several future directions arising from this work.
Evidently, ancilla systems are particularly affordable when only one layer of additional qubits is necessary. We were able to find logical operators on the gross code where this is the case. Are there any other families of codes where this is possible? Our analytical techniques did not suffice to prove that a monolayer system suffices for the gross code. Is it possible to prove that a single layer is enough? Quantum expander codes seem promising candidates.
For our simulations, we derived explicit syndrome circuits for the $\bar{X}$ and $\bar{Z}$ measurement schemes for the gross code. Are these syndrome circuits optimal in terms of the number of fault locations? Can one design optimal or otherwise high-performance syndrome circuits for other measurement schemes, perhaps assuming that the original code $\mathcal{G}$ also has an efficient syndrome circuit?
Finally, from Section 4.3 it is clear that implementing logical computations will require very many measurements, which in turn required several rounds of error correction. For certain error correction codes, such as perhaps [SHR24], is there a single-shot error correction scheme for merged codes? Lowering the time overhead of surgery operations would have significant practical and theoretical implications.
@@ A.1 Proofs for the $X$ Ancilla System
Finally, we show that if $\mathcal{G}$ has $k$ logical qubits, then $\mathcal{G}_{X}$ has k1k-1 logical qubits. We proceed by induction on the number of layers, and we use $\mathcal{G}^{l}_{X}$ to indicate the merged code for measuring $\bar{X}_{M}$ with $l$ layers, and $H^{X}_{l}$, $H^{Z}_{l}$ its parity check matrices.
@@
The size of $U_{1}$ is the rank of $G$. Observe that the rank of $F$ is |V|1|V|-1, since we assumed that the only logical operator supported on $V$ is $\bar{X}_{M}$. Therefore $|U_{1}|=\operatorname{rank}(G)=|C_{0}|-\operatorname{rank}(F)=|C_{0}|-|V_{0}|+1$, and the number logical qubits in $\mathcal{G}_{X}^{1}$ is
@@
We proceed with the inductive step: assuming that $\mathcal{G}^{L-2}_{X}$ has k1k-1 logical qubits, we show the same for $\mathcal{G}^{L}_{X}$. The newly added set of $X$ checks $V_{L}$ is linearly independent. A product of any subset of $V_{L}$ will have non-trivial support on $C_{L}$, and while $H_{X}(V_{L})=X(V_{L-1})$ has no support on $C_{L}$, by assumption it cannot be a stabilizer in $H^{X}_{L-2}$. Similarly, the newly added $Z$ checks $C_{L-1}$ is linearly independent, and their products always have support on $C_{L}$. We conclude that $\mathcal{G}_{X}^{L}$ has k1k-1 logical qubits.
In fact, these k1k-1 qubits are the remaining logical qubits from $\mathcal{G}$. For any $\bar{Z}$ that commutes with $\bar{X}_{M}$, its support on $V_{0}$ is even in size, and as a consequence of Lemma 8, we can clean its support from $V_{0}$ by multiplying by $Z$ checks of $\mathcal{G}$. This implies some $\bar{Z}^{\prime}\equiv_{\mathcal{G}}\bar{Z}$ supported only on $R$, which is then also a logical operator of $\mathcal{G}_{X}$. For a logical operator $\bar{X}\not\equiv_{\mathcal{G}}\bar{X}_{M}$ of $\mathcal{G}$ supported only on $V_{0}\cup R$, notice that $\bar{X}$ is also a logical operator of $\mathcal{G}_{X}$. It is nontrivial in $\mathcal{G}_{X}$ because there was some $\bar{Z}$ that anticommuted with $\bar{X}$ in $\mathcal{G}$ and commuted with $\bar{X}_{M}$, and therefore the construction of $\bar{Z}^{\prime}$ still anticommutes with $\bar{X}$ in $\mathcal{G}_{X}$. ∎
@@ A.2 Proofs for the $Y$ Ancilla System
which implies the number of logical qubits is k1k-1.
Inductively, we make the same observation as in the proof of Theorem 1: the newly added set of $X$ checks $V^{X}_{L}$ is linearly independent. Any $H_{X}(v\in V^{X}_{L})$ will have non-trivial support on $C_{L}$, and while $H_{X}(V^{X}_{L})=X(V_{L-1})$ has no support on $C_{L}$, by assumption it cannot be a stabilizer in $\mathcal{G}^{L-2}_{Y}$. Similarly, the newly added $Z$ checks in $C^{X}_{L-1}$ are linearly independent, and their products always have support on $C_{L}$. We conclude that $\mathcal{G}^{L}_{Y}$ has k1k-1 logical qubits, which are the original logical qubits from $\mathcal{G}$. ∎
@@ A.3 Proofs for the Multi-block $XX$ Ancilla System
[…]
@@
$\times\left(\prod_{\begin{subarray}{c}j=1\\ j\text{ odd}\end{subarray}}^{L}X(v_{j}F^{(1)\top}\in C_{j}^{(1)})\right).$ (101)
Since $|X(B)S_{1}|$ is at least the weight of factors in line 100, we proceed by ignoring the factors in line 101.